Air at 10.0◦C and normal atmospheric pressure is pumped into a car tire, which causes the air to compress to 28.0% of its original volume and heat up to 40.0◦C

(a) What is the tire pressure?

(b) You take the car for a drive at high speeds, and the air temperature in the tire rises to 85.0◦C while the interior volume increases by 2.00%.   What is the new absolute

pressure (in pascals)?
Answer #1

a) using the equation

P1V1 / T1 = P2V2 / T2

1*V / (273 + 10) = P2*0.28V / (273 + 40)

P2 = 3.95 atm

so the tire pressure is 3.95 atm

Tire pressure in pascal = 3.95*1.01325*10^5 = 4*10^5 Pa

(b) 4*10^5 * V / (273 + 40) = P2 * 1.02V / (273 + 85)

P2 = 4.488*10^5 Pa